Consider a put option on a share of Apple Stock (current spot price = $175) with an expiration of

March 15, 2019 and a strike of $170. One contract is for 50 shares and has a premium of $450.00.

What is the breakeven point of this contract?

If the spot price of Apple at maturity is $165, then what is the profit from one long contract?
Answer #1

Let Breakeven Point is when share price on the expiration date is $x
Therefore, (170-x)*50 = 450
which gives us , x = 170 – 150 = 170 – 9 = 161

Hence, breakeven point of this contract is share price at the expiration date = $161

If the spot price of Apple at maturity is $165, then what is the profit from one long contract = 50*($170-$165) - $450 = -$200
Hence, there will be a loss of $200

Please do rate me and mention doubts, if any, in the comments section.